# Integration of Beyond Diagonal RIS and UAVs in 6G NTNs: Enhancing Aerial Connectivity

### v1

@@ I Introduction
Currently, research on UAV networks predominantly emphasizes conventional RIS, specifically diagonal RIS (D-RIS), which employs diagonal phase shift matrices to manipulate incident signals. In D-RIS, each element independently adjusts the wireless environment without coordination among elements [5]. While D-RIS offers a simple design that effectively enhances signal strength, its limited inter-element coordination restricts its beamforming capabilities, thereby reducing its overall performance in more complex communication scenarios This article investigates the advantages of BD-RIS when considered in combination with UAVs, which enables enhanced control over the wireless environment by facilitating connections among reconfiguring elements [6]. BD-RIS not only accomplishes a complete scattering matrix, but it also provides full space coverage by incorporating signal reflection and refraction [7]. This cutting-edge concept has the capacity to completely transform the efficiency and adaptability of UAV networks.
This article provides the integration of BD-RIS and UAV communications in 6G NTNs. First, we discuss the preliminaries of UAV communications, including applications and use cases of UAV communications in Section II. Then, in Section III, we present the fundamentals of BD-RIS, discuss the difference between BD-RIS and D-RIS and their key characteristics, analyze the hardware architectures of BD-RIS, and propose various configurations of UAV-based BD-RIS systems. Subsequently, in Section IV, we provide the key potentials when integrating BD-RIS and UAV communications in 6G NTNs. We also provide a technical case study based on transmissive BD-RIS mounted UAV communications and the results are compared with the benchmark transmissive D-RIS mounted UAV communications in Section V. It is shown that BD-RIS outperforms D-RIS in improving the spectral efficiency of UAV communications. Finally, in Section VI, we highlight future research directions with a concluding discussion in Section VII.
@@ II-B3 Relay Node
Comparison Diagonal IRS Beyond Diagonal IRS
@@ IV-D Strengthened Security and Privacy
Security is a big issue since UAV communications are prone to cybersecurity problems like possible hacking and data leaks. The protection of private data and guarantee of operational integrity depend on strong encryption and safe communication methods being used. Privacy problems also surface primarily in relation to surveillance or monitoring UAVs are employed for. BD-RIS can improve the security and privacy of UAVs by helping to maximize signal pathways, which helps to lower the possibility of eavesdropping or interception. Furthermore, the BD-IRS can be made to incorporate security elements safeguarding data flow between UAVs and the network, therefore addressing possible cybersecurity risks and safeguarding privacy.
@@ VI-2 Cooperative Control of UAV-BD-RIS
Another potential research direction is optimizing both the configurations of BD-RIS and the trajectories of UAVs together. Simultaneously optimizing the settings of the RIS and the trajectories of UAVs can significantly enhance communication performance. Implementing real-time adaption algorithms for UAV placements and IRS positions, based on instantaneous channel information and user mobility patterns, can significantly improve the responsiveness and reliability of the system.

### v2

@@ I Introduction
Currently, research on UAV networks predominantly emphasizes conventional RIS, specifically diagonal RIS (D-RIS), which employs diagonal phase shift matrices to manipulate incident signals. In D-RIS, each element independently adjusts the wireless environment without coordination among elements [5]. While D-RIS offers a simple design that effectively enhances signal strength, its limited inter-element coordination restricts its beamforming capabilities, thereby reducing its overall performance in more complex communication scenarios. The concept of Beyond Diagonal Reconfigurable Intelligent Surfaces (BD-RIS) has recently been introduced, extending the functionality of reflecting surfaces by enabling a full scattering matrix that is not restricted to a diagonal structure [6]. This advanced capability allows BD-RIS to perform sophisticated beamforming, providing simultaneous support for multiple users and facilitating complex signal manipulation. This article explores the advantages of integrating BD-RIS with UAVs, offering enhanced control over the wireless environment through element interconnection [7]. In addition to achieving a complete scattering matrix, BD-RIS enables full space coverage by incorporating both signal reflection and refraction [8]. This innovative concept has the potential to revolutionize the efficiency and adaptability of UAV networks. To the best of our knowledge, this is the first work to explore the integration of BD-RIS and UAV communication in 6G NTNs.
The main contributions of this article can be summarized as follows. First, we discuss the preliminaries of UAV communications, including applications and use cases of UAV communications in Section II. Then, in Section III, we present the fundamentals of BD-RIS, discuss the difference between BD-RIS and D-RIS and their key characteristics, analyze the hardware architectures of BD-RIS, and propose various configurations of UAV-based BD-RIS systems. Subsequently, in Section IV, we provide the key potentials when integrating BD-RIS and UAV communications in 6G NTNs. We also provide a technical case study based on transmissive BD-RIS mounted UAV communications and the results are compared with the benchmark transmissive D-RIS mounted UAV communications in Section V. It is shown that BD-RIS outperforms D-RIS in improving the spectral efficiency of multi-user UAV communications. Finally, in Section VI, we highlight future research directions with a concluding discussion in Section VII.
@@ II-B Use Cases of UAV Communications
As shown in Fig. 1, a UAV can serve as a flying UE, an aerial BS, or a flying relay.
@@ II-B3 Flying Relay
Comparison Diagonal RIS Beyond Diagonal RIS
@@ III-B3 Group-connected BD-RIS
III-B4 Architectural Trade-off
The best-performing BD-RIS architecture is the fully-connected architecture, as it provides advanced beamforming capabilities, supports multiple users simultaneously, and achieves high spectral and energy efficiency. However, it has the highest complexity due to intricate interconnections and control requirements. On the other hand, the single-connected architecture is the least complex but offers limited performance. The group-connected architecture balances performance and complexity, making it a practical choice for many applications. In addition, BD-RIS can function as a transmitter by actively modulating signals, as a reflector/refractor to enhance coverage and mitigate blockages, or, theoretically, as a receiver by processing and absorbing incoming signals. While BD-RIS is predominantly used as a transmitter or reflector/refractor in communications, its application as a receiver is less explored and might be relevant in advanced system designs like BD-RIS-assisted MIMO networks.
@@ IV-D Strengthened Security and Privacy
Security is a big issue since UAV communications are prone to cybersecurity problems like possible hacking and data leaks. The protection of private data and guarantee of operational integrity depend on strong encryption and safe communication methods being used. Privacy problems also surface primarily in relation to surveillance or monitoring UAVs are employed for. BD-RIS can improve the security and privacy of UAVs by helping to maximize signal pathways, which helps to lower the possibility of eavesdropping or interception. Furthermore, the BD-RIS can be made to incorporate security elements safeguarding data flow between UAVs and the network, therefore addressing possible cybersecurity risks and safeguarding privacy.
@@ VI-2 Cooperative Control of UAV-BD-RIS
Another potential research direction is optimizing both the configurations of BD-RIS and the trajectories of UAVs together. Simultaneously optimizing the settings of the RIS and the trajectories of UAVs can significantly enhance communication performance. Implementing real-time adaption algorithms for UAV placements and RIS positions, based on instantaneous channel information and user mobility patterns, can significantly improve the responsiveness and reliability of the system.
@@ VII Conclusion
Biographies
Wali Ullah Khan [M] (waliullah.khan@uni.lu) received a Ph.D. degree in information and communication engineering from Shandong University, China, in 2020. He is currently working with the SIGCOM Research Group, SnT, University of Luxembourg. His research interests include 6G wireless communications, satellite communications, integrated terrestrial and non-terrestrial networks.
Eva Lagunas [SM] (eva.lagunas@uni.lu) received a Ph.D. degree in telecommunications engineering from the Polytechnic University of Catalonia (UPC), Barcelona, Spain, in 2014. She currently holds a research scientist position in the SIGCOM Research Group, SnT, University of Luxembourg.
Asad Mahmood [S] (asad.mahmood@uni.lu) received his Master degrees in Electrical Engineering from COMSATS University Islamabad, Wah Campus, Pakistan, in 2020, and PhD degree in Computer Science from University of Luxembourg, Luxembourg in 2025. He is currently working as a Research Associate with the SIGCOM Research Group, SnT, University of Luxembourg.
Muhammad Asif (masif@ujs.eu.cn) received a Ph.D. degree in information and communication engineering from the University of Science and Technology of China (USTC), Hefei, China, in 2019. He is currently working as a post-doctoral researcher at the School of Computer Science and Communication Engineering, Jiangsu University, Zhenjiang, China. His research interests include 6G wireless communications, and integrated terrestrial and non-terrestrial wireless networks.
Manzoor Ahmed (manzoor.achakzai@gmail.com) received a Ph.D. from Beijing University of Posts and Telecommunications China in 2015. He is currently a professor with the School of Computer and Information Science and also with the Institute for AI Industrial Technology Research, Hubei Engineering University, Xiaogan, China.
Symeon Chatzinotas [F] (symeon.chatzinotas@uni.lu) received Ph.D. degree in electronic engineering from the University of Surrey, Guildford, United Kingdom, in 2009. He is currently a full professor and head of the SIGCOM Research Group, SnT, University of Luxembourg.